OpenStudy (anonymous):

How do I work out the problem 9^x-3^x-12=0 ? logarithmic function

OpenStudy (anonymous):

gimmick is to write \[9^x\] as \[3^{2x}\] and solve a quadratic equation

OpenStudy (anonymous):

i know the answer but trying to get teh answer is my problem

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im confused because you are using a different example :-(

OpenStudy (anonymous):

original equation is \[9^x-3^x-12=0\]

OpenStudy (anonymous):

put \[z=3^x\] get \[z^2-z-12=0\]

OpenStudy (anonymous):

factor as \[(z-4)(z+3)=0\] \[z=4, z=-3\]

OpenStudy (anonymous):

so \[3^x=4\] or \[3^x=-3\] and the second one is not possible so only \[3^x=4\]

OpenStudy (anonymous):

that's wrong it can't be two different answers though?

OpenStudy (anonymous):

now solve for x. one solution is \[x=\log_3(4)\]

OpenStudy (anonymous):

no only one answer

OpenStudy (anonymous):

the quadratic has two answers, but \[3^x>0\] for all x so you can ignore \[3^x=-3\]

OpenStudy (anonymous):

i'm all ears

OpenStudy (anonymous):

\[3^x=4\]

OpenStudy (anonymous):

i used a calculator and got -12 as an answer

OpenStudy (anonymous):

you want an actual number out of this right?

OpenStudy (anonymous):

it says to solve the equation: 9 to the power of x minus 3 to the poewr of x minus 12 = 0

OpenStudy (anonymous):

\[3^x=4\iff x = log_3(4)\iff x = \frac{\ln(4)}{\ln(3)}\]

OpenStudy (anonymous):

got it, and you solution is above

OpenStudy (anonymous):

if you want a decimal use a calculator,

OpenStudy (anonymous):

i need to read it again. yes i can figure it out with a calculator :-) one sec

OpenStudy (anonymous):

i will get rid of the junk

OpenStudy (anonymous):

okay i prefer that :D

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

how'd it go?

OpenStudy (anonymous):

im just confused can you put exactly what u said in one line? :P so i can read it

OpenStudy (anonymous):

to work it out, not the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the trick is to recognize \[9=3^2\] so \[9^x=(3^2)^x=3^{2x}\] then you have \[3^{2x}-3^x-12=0\] which is a quadratic equation in \[3^x\] because \[3^{2x}=(3^x)^2\] so to make life easy replace \[z=3^x\] giving you the equation \[z^2-z-12=0\] \[(z-4)(z+3)=0\] \[z=4,z=-3\] now go back and replace \[z\text{ by }3^x\] and get \[3^x=4\] or \[x^x=-3\] but there is no not about those damned way that \[x^x=-3\] so only \[3^x=4\] is a solution now solve for x. you could say \[x=\log_3(4)\] but that doesn't really give you a number for x. if you want a decimal use the change of base formula to get \[x=\frac{\ln(4)}{\ln(3)}\] and then a calculator

OpenStudy (anonymous):

wow i don't know where that line came from. what i meant is THERE IS NO WAY THAT \[3^x=-3\]

OpenStudy (anonymous):

ok i understand, answer has to be positive.

OpenStudy (anonymous):

But when I input 4 for x , 9^4-3^4-12 doesn't equal zero?

OpenStudy (anonymous):

woops ignore that last msg

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

you are right I get it now. Answer is correct.

OpenStudy (anonymous):

hurrah!

OpenStudy (anonymous):

this is my first time here, how do i give you a medal?

OpenStudy (anonymous):

i figured that out :P thanks so much.